how do I find the Focus, Equation of the Directrix, and the equation of the axis of symmetry from the equation
x=(1/16)(y-7)^2-5

Answer by stanbon(75887)

how do I find the Focus, Equation of the Directrix, and the equation of the axis of symmetry from the equation
x=(1/16)(y-7)^2-5
------
Since the vertex is at (-5,7),
axis of symmetry: y = 7
Focus and Directrix ????
4p = 1/16
p = 1/64
----
Focus: (-5+(1/64),7)
Directix: x = -5-(1/64)
=============================
Cheers,
Stan H.
================